The equation of a line paralell to the given line but passing through the given point
Y= 1/4x-2; (8,-1)

Answer by Alan3354(69443)

The equation of a line paralell to the given line but passing through the given point
Y= 1/4x-2; (8,-1)
-------------
Slope m = 1/4
---
--> y+1 = (1/4)*(x-8)